randomly selected weights in ounces represent some of the a biological study on a terret breeding population in Wyoming. 1. The following 5, 9, 6, 6, 9, 11, 8, 3, 5, 5, 5, 9, 7, 9, 11, 11, 23, 5, 8, 9, 6, 8, 7 data type of the measure compute the five number summary summary figures does this dataset have any outliers? (a) Define the population, element of study and the measure. Discuss the (b) Sort the dataset, construct the Data Table, draw the histogram and (e) calculate and fully define the Interquartile Range. Using these 2. The following closing prices were recorded from a sample of randomly selected commodities listed on the Chicago Mercantile Bxchange. Each value represents the commoditys value at the close of trading on January 31, 2014. 132, 125, 122, 129, 128, 125, 122, 120, 123, 139, 136, 130, 120, 122, 122 (a) Define the population, unit of sample, the measure for these (b) construct the date table, graph the relative frequencies and compute c) Calculate and fully define the standard deviation. Using these summary observations and fully define the Range the mean and variance. figures what is the most common closing prices for all other commodities? 3. A random experiment is conducted by flipping a Quarter five times and recording the number of the heads that appear on the top side. List some of the samp the number of heads in five flips. What do we mean by the term Sample Space? le points in the space and construct data table for each possible value of a) Draw the relative frequency distribution and list all the probabilities heads for the statistic, the number of heads, or each possible count of b) What is the average number of heads you would expect to see if you conducted the experiment many times? What is the number of heads you would expect to see in the experiment for any one repetition of the experiment? probability in your own words and explain why such a calculation is important in statistical analysis c) What is the probability that I flip two or more heads? Define

Hi I would love some help on these problems. It is mainly to ensure I followed the correct steps to solving these problems including all graphs and tables. thank you in advance.
Answer #1

1.

a) Population: Weights in ources on a ferre breeding populationin wyoming
These measures are weights
b) Data Display

Sort
3 5 5 5 5 5 6 6 6 7 7 8 8 8 9
9 9 9 9 11 11 11 23

The frequency distribution is

LL UL f
3 6.9 9
7 10.9 10
11 14.9 3
15 18.9 0
19 22.9 1
Total = 23

Histogram 12 10 10 9 4 0 0 3-7 7-11 11-15 15-19 19-23

Minmann value = 3 First Quarile 2341) Rank IInd Quarile 2-Rank 234 บั 12thrank 2 ThirdQuartile Rank 23+1 Rank =18 Rank = 9 Ma

c)

IQR = Q3-Q1 = 9-5-4

There is one Outlier
Number of data values below Q1 by more than 1.5 IQR: 0
Number of data values above Q3 by more than 1.5 IQR: 1 i..e. 23